OpenStudy (anonymous):

7/8p-4=10 I need the value of P

OpenStudy (anonymous):

Given:\[\frac{ 7 }{ 8p }-4=10\] Add four, multiply each side by (8p)/1 and you should be able to simplify from there.

OpenStudy (anonymous):

Rearrange the equation to equal P 7/8p-4=10 7/8p=14 (plus 4 to both sides) 8p=0.5 p=1/16 (divide both sides by 8 to find p)